OpenStudy (anonymous):

Which proportion models this problem? On a soccer team, 11 out of 17 players surveyed say they had two or more siblings. The league has 850 players. Which is the best prediction of the number of players in the league that have two or more siblings? A. B. C. D.

OpenStudy (anonymous):

A. x/11=17/850 B. x/17=11/850 C. x/850=17/11 D. x/850=11/17